OpenStudy (anonymous):

4 x 5^3.2a = 29 Salve each equation

OpenStudy (anonymous):

\[ \begin{align} 4 \cdot 5^{3.2a} &= 29\\ 5^{3.2a} &= \frac{29}{4}\\ \ln{5^{3.2a}} &= \ln{\frac{29}{4}}\\ 3.2a\cdot \ln{5} &= \ln{\frac{29}{4}}\\ a &= \frac{\ln{29/4}}{3.2\ln5} \end{align} \]

OpenStudy (anonymous):

Where did ln5 go?

OpenStudy (anonymous):

Divided through.